An exponential is such that the 3rd term minus the first term is 48. The fourth term minus the second term is 44. Find the common ratio and the first term

Answer by richwmiller(17219)

Tn = t * r^(n - 1)
T3 = t * r^(2)
T1 = t
t * r^2-t=48
T4 = t * r^3
T2 = t * r
t*r^3- t*r=44,
t*r^2-t=48
r = 11/12, t = -6912/23
I suspect there is a mistake since this is similar with 44 instead of 144